OpenStudy (ericanoel912):

FAN AND MEDAL! If f(x) = log2 (x + 4), what is f−1(3)?

OpenStudy (anonymous):

You need to find the inverse of log 2 (x+4)....

OpenStudy (ericanoel912):

Okay

OpenStudy (ericanoel912):

How do I do that? @nelsonjedi

OpenStudy (anonymous):

So if we make f(x) = y what would our inverse equation look like?

OpenStudy (ericanoel912):

Um... I know i have to set it equal to y and then switch x and y, then solve for y.... but I don't know how because of the "log" part.

OpenStudy (anonymous):

Inverse: Step one make f(x) = y so our equation now looks like y= log2(x+4)..correct?

OpenStudy (ericanoel912):

Yes

OpenStudy (anonymous):

Right..now switch the x and y and what will it look like?

OpenStudy (ericanoel912):

then x=log2(y+4) ?

OpenStudy (ericanoel912):

then x-4/log2=y ???

OpenStudy (ericanoel912):

??

OpenStudy (anonymous):

Hold on..

OpenStudy (ericanoel912):

Okay

OpenStudy (anonymous):

Now you need to convert the equation from a logarithmic equation to an exponential equation

OpenStudy (ericanoel912):

Uh-oh..

OpenStudy (ericanoel912):

I really have to finish this... almost dinner. lol

OpenStudy (anonymous):